# From Learning to Safety: A Direct Data-Driven Framework for Constrained Control

### v1

@@ 7 Comparison with other approaches and limitations
Comparison with safe RL: Existing methods for safe RL include using reward shaping [25, 7], Lagrangian methods combined with policy gradient or actor-critic algorithms [43], interior-point optimization [44], and safety filtering [18, 45]. Additionally, a small body of work explores direct stochastic optimization of neural network controllers over finite horizons, improving safety at sampled states but demanding significant computational resources [46]. However, as discussed in the introduction, because learning algorithms operate stochastically, learning-based control—particularly when an explicit model is unavailable—requires safety filters to regulate policy execution and to ensure reliable safety. The proposed optimization-based control framework is compatible with all the safe RL methods mentioned above, and provides formal safety assurance under learning errors.
Comparison with our previous work [23]: In [23], the definition of SACBFs is proposed, which contributes to a safety enhancement framework analogous to (3). The current work expands [23] in the following two aspects. First, all the learning-based methods for SACBFs proposed in Section 4 totally remove the assumption on the availability of a nominal model, which is required in [23]. Second, in [23], the robustness to learning errors is addressed by a new CBF with a contractive safe set, which needs a stronger invariance condition than invariant sets. In contrast, the current work establishes an ESSf analysis framework using the fundamental robustness property of the SACBF, resulting in significantly less conservative constraint tightening than [23].
@@
We extend the barrier certificate, originally used to characterize state safety, to the SACBF, which captures both state-input safety. However, this extension introduces a drawback: it necessitates sampling and learning in the state-input space rather than just the state space, thereby increasing computational and sample complexity.
@@ 8.1 Model
where $L=0.1$. The state vector $x$ includes the position $p_{x}$, $p_{y}$, the speed $v$, and the yaw angle $\Psi$. The acceleration $a$ and the steering angle $\delta$ are the inputs. The input constraints are given by $-5\leq a\leq 2$ and $|\delta|\leq\pi/4$. The state constraints are specified by $|p_{x}|\leq 2.6$, $|p_{y}|\leq 2.6$, $0\leq v\leq 1$, and $|\Psi|\leq\pi$, as well as the requirement of avoiding some obstacles shown in Fig. 2. According to the considered state constraints, the function $h$ is as follows:
@@ 8.2 Synthesizing SACBFs
The synthesis of the SACBF is performed using three approaches, including the expert-guided learning approach presented in Section 4.2, the RL approach described in Section 4.3, and the LMI approach of Section 4.4. The SL approach in Section 4.1 is not considered because it is unrealistic to assume knowing a CBF without knowing the model in this example. For the expert-guided approach, we adopt artificial potential fields (APF) [47] to design the expert controller. The APF-based controller has a certain capability for obstacle avoidance but is prone to falling into local optima, which can either reside within obstacle regions or be located far from the target.
@@
Figure 3: The boundaries (blue curves) of the safe set $\{x|x^{T}P^{-1}x\leq 1\}$ of the quadratic SACBF (4.4) obtained by solving LMIs
For the LMI approach, since Assumption 3 cannot be satisfied for the original vehicle model in (37), we make some simplifications. In particular, (i) only the subsystems of $p_{y}$ and $\Psi$ are considered, (ii) the speed is fixed to a constant $0.5$, (iii) the obstacles are removed, and (iv) the approximation $\tan\delta\approx\delta$ is used. Then, the resulting simplified system is of the form (19), with the nonlinear term $\sin\Psi$. After using the input signal $\delta(t)=0.1\sin t$ to excite the system, we get the state-input data $\mathbf{D}$ (defined in Lemma 2) with the time horizon $20$. The persistency of excitation condition is satisfied. By solving the LMI problem (27), we obtain $Q^{B}_{P}$ with $P=\left[\begin{array}[]{cc}6.25&-0.2996\\ -0.2996&0.1176\end{array}\right]$. The corresponding safe set is visualized in Fig. 3. As observed, the LMI approach is conservative, as it restricts the SACBF candidate to quadratic forms.
@@ 8.3 Closed-loop simulation
Fig. 4(a) demonstrates the performance of the 3DSF learned using the expert-guided learning approach. When applying the unsafe policy, the trajectories exhibit some constraint violations. In contrast, the learned SACBF enables the vehicle to smoothly avoid obstacles, except for the initial condition $p_{x}=2,\;p_{y}=0.5$ (the rightmost black dot). To explain this constraint violation, we note that the expert controller (safe APF) also exhibits constraint violations when starting from this initial state. Consequently, the 3DSF, which is learned based on this expert controller, recognizes this initial state as unsafe and therefore fails to refine the reference policy.
@@
In comparison, the 3DSF learned by RL refines the reference policy for all the initial states. The trajectory starting from the rightmost initial state in Fig. 4(b) demonstrates the superiority of the proposed 3DSF learned by RL. In particular, the safe APF policy makes the trajectory fail into a local optimum inside the right obstacle, while the unsafe APF policy with the 3DSF circumvents the local optimum and steers the system to the target.
@@
In Figs. 4(c-d), we show the trajectories of the vehicle controlled by the DDPG policies. It is found that although for some initial states safe DDPG successfully plans a safe trajectory, some trajectories fail to converge to the target set. This illustrates that it is hard to balance the task performance and constraint satisfaction by naively adding penalties to the stage cost during training. In contrast, the safety filter significantly enhances the safety performance of the unsafe DDPG policy while ensuring that the policy reaches the target.
@@
(a) Closed-loop trajectories. Dashed blue curve: unsafe APF; Solid blue curve: unsafe APF with 3DSF learned from the expert controller; Red curve: safe APF; Black dots: initial states.
@@
(b) Closed-loop trajectories. Dashed blue curve: unsafe APF; Solid blue curve: unsafe APF with 3DSF learned by RL; Red curve: safe APF; Black dots: initial states.
@@
(c) Closed-loop trajectories. Dashed blue curve: unsafe DDPG; Solid blue curve: unsafe DDPG with 3DSF learned from the expert controller; Red curve: safe DDPG; Black dots: initial states.
@@
(d) Closed-loop trajectories. Dashed blue curve: unsafe DDPG; Solid blue curve: unsafe DDPG with 3DSF learned by RL; Red curve: safe DDPG; Black dots: initial states.
@@
Figure 4: Comparison of closed-loop vehicle trajectories under different controllers and different safety filters.
@@ 8.4 Statistical evaluation and comparison with indirect data-driven safety filters
Figure 5: Comparison of closed-loop vehicle trajectories under the 3DSF and the indirect safety filter.
Finally, we statistically compare the proposed 3DSF against an indirect data-driven safety filter. In this subsection, the SACBF is learned by the RL approach because it has been illustrated in the previous subsection that the SACBF learned by the RL approach results in a larger safe set. To design the indirect data-driven safety filter, we intend to learn a standard CBF based on an approximate model. We first use a neural network with [128 128 128] “elu” hidden layers to identify the kinematic model. After 30 epochs of training, we get an approximate model with a sufficiently small ($2.38\cdot 10^{-4}$) mean square error. Then, following the approaches of [12, 8, 23], the approximate model is used as a prediction model in the reachability problem [^4] The reachability problem (4.3) is defined over an infinite horizon. To make it solvable for each sampled state, we truncate the horizon at 20. While a longer horizon provides a better approximation of the maximal safe set, a horizon of 20 balances between sample complexity and safety performance. (4.3). We solve (4.3) with the initial state $x_{0}$ being each state sample obtained in Section VIII.B. Then, we get the target value for the neural CBF, which is parameterized by a neural network with [128 128 32] “tansig” hidden layers.
Table 1: Comparison of safety rate, total cost, and average online CPU time for different control policies and safety filters. The total cost is defined as the sum of $p^{2}_{x}+p^{2}_{y}$ over 1000 time steps. The safety rate is defined as the ratio between the number of successful trajectory plans (without constraint violations) and the number of all trajectory plans. “SF” means “safety filter”.
@@
The results in Table 1 provide a detailed comparison of the safety rate, total cost, and average CPU time for those trajectories starting from the sampled initial states. The total cost reflects the ability of governing the vehicle to the target set. The key findings are summarized below:
@@ 9 Conclusions and future work
We have proposed an optimization-based control framework that contains an SACBF constraint to ensure safety for learning-based control methods. The main advantages lie in its universal applicability to RL and SL methods for designing controllers, as well as its online computational efficiency. Three learning algorithms have been developed within the optimization-based control framework, making the controller strive for optimal performance while ensuring safety to the greatest extent possible. We have analyzed the theoretical properties regarding the robustness of SACBF and translated the results to error-to-state safety (ESSf) of the proposed control framework w.r.t. learning errors. Simulations conducted on an obstacle avoidance problem demonstrate the aforementioned advantages.
Future work will focus on (i) increasing the scalability to higher dimensional systems, (ii) examining the effect of external uncertainties, and (iii) designing distributed optimization-based control for multi-agent systems subject to unknown dynamics and nonlinear constraints.

### v2

@@ 7 Discussions
The three approaches we proposed can be viewed as extensions of traditional CBF synthesis methods (typically model-based) [44, 49, 12] to the synthesis of SACBFs in a model-free framework. Since SACBFs remove the model dependency by enforcing the invariance condition directly on the joint state-action space, learning SACBFs requires learning a safety function over a higher-dimensional domain with dramatically more samples and without relying on an explicit model. Secondly, since the safe set is not explicitly defined by the level set, SACBF learning must jointly infer the safe set, the SACBF function, and the invariance structure. Finally, ESSf robustness analysis is more involved for SACBFs because it must characterize the effect of learning errors without propagating them through the system model. For these reasons, learning SACBFs involves fundamentally different technical challenges compared to learning CBFs.
@@
Comparison with safe RL: Existing methods for safe RL include using reward shaping [35, 23], Lagrangian methods combined with policy gradient or actor-critic algorithms [54], interior-point optimization [34], and safety filtering [10, 13]. Additionally, a small body of work explores direct stochastic optimization of neural network controllers over finite horizons, improving safety at sampled states but demanding significant computational resources [31]. The work of [18] approximates the optimal safe policy by using RL to solve a time-discounted modification of the Hamilton-Jacobi reachability problem, while [27] extends this framework to adversarial RL, making the resulting policy robust to model uncertainty. More recently, [30] establishes the Lipschitz continuity of the discounted value function introduced in [18] and leverages this property to develop certification methods for the corresponding approximate policy. However, as discussed in the introduction, because learning algorithms operate stochastically, learning-based control—particularly when an explicit model is unavailable—requires safety filters to regulate policy execution and to ensure reliable safety. The proposed optimization-based control framework is compatible with all the safe RL methods mentioned above, and provides formal safety assurance under learning errors.
Comparison with existing work on SACBF-based safety: In [24], a general definition of SACBFs is introduced. In parallel, the authors of [18] and [40] study the synthesis of SACBFs through model-free learning. The current work extends [24, 18, 40] in three major ways. First, we propose three learning-based methods for synthesizing SACBFs under different assumptions on the prior knowledge about safety information, without requiring any nominal model, which is assumed in [24]. That means that the SACBF considered in this paper is not limited to the state-action value functions in [24, 18, 40]. Second, to address the effect of learning errors, this paper establishes robustness analysis via the notion of ESSf, resulting in significantly less conservative constraint tightening than [24]. In contrast, [18] and [40] do not examine the effect of learning errors. Third, unlike the safety-filter approaches in [24, 18, 40], our optimization-based controller uses a general nonlinear objective $Q$ and a policy refinement strategy to achieve near-optimal performance while preserving safety.
@@
In the paper, we have extended the barrier certificate, originally used to characterize state safety, to the SACBF, which captures both state-input safety. However, this extension introduces a drawback: it necessitates sampling and learning in the state-input space rather than just the state space, thereby increasing computational and sample complexity. In general, the sample complexity for providing formal performance guarantees (see Section V) is exponential with respect to the dimension of the state-input space. This is also the case for other methods from the literature [44, 38, 25, 23]. Potential solutions to mitigate this complexity include leveraging model symmetries or reduced-order representations, which will be investigated in our future work.
We focus on discrete-time nonlinear systems because our work targets safety in learning-based control methods, like RL and SL, which typically use discrete-time models. The extension of SACBFs and the associated robustness analysis to continuous-time dynamical systems will be investigated in our future work.
@@ 8.1 Model
where $L=0.1$. The state vector $x$ includes the position $p_{x}$, $p_{y}$, the speed $v$, and the yaw angle $\Psi$. The acceleration $a$ and the steering angle $\delta$ are the inputs. The input constraints are given by $5\leq a\leq 2$ and $|\delta|\leq\pi/4$. The state constraints are specified by $|p_{x}|\leq 2.6$, $|p_{y}|\leq 2.6$, $0\leq v\leq 1$, and $|\Psi|\leq\pi$, as well as the requirement of avoiding some obstacles shown in Fig. 2. According to the considered state constraints, the function $h$ is as follows:
@@ 8.2 Synthesizing SACBFs
The synthesis of the SACBF is performed using two approaches, including the expert-guided learning approach presented in Section 4.2 and the RL approach described in Section 4.1. The SL approach in Section 4.3 is not considered because it is unrealistic to assume knowing a CBF without knowing the model in this example. For the expert-guided approach, we adopt artificial potential fields (APF) [51] to design the expert controller. The APF-based controller has a certain capability for obstacle avoidance but is prone to falling into local optima, which can either reside within obstacle regions or be located far from the target.
@@
Fig. 3 demonstrates how the size of $S_{\omega}$ and $X_{\omega}$ varies with different values of $\beta$ when applying the expert-guiaded learning approach. From Fig. 3, we can observe that the safe set is non-empty for all $\beta=0,\;0.2,\;0.4,\;0.6,\;0.8$.
Figure 3: Illustration of the safe set learned from the expert controller under different $\beta$.
@@ 8.3 Closed-loop simulation
Fig. 4(a) demonstrates the performance of the 3DSF learned using the expert-guided learning approach. When applying the unsafe policy, the trajectories exhibit some constraint violations. In contrast, the learned SACBF enables the vehicle to smoothly avoid obstacles, except for the initial condition $p_{x}=2,\;p_{y}=0.5$ (the rightmost black dot). To explain this constraint violation, we note that the expert controller (the safe APF) also violates the constraints when initialized from this state. This occurs because the APF becomes trapped in a local optimum located inside the obstacle in the right half-plane. Consequently, the 3DSF, which is learned based on this expert controller, recognizes this initial state as unsafe and therefore fails to refine the reference policy. The green curve in Figs. trajectory(a) and trajectory(c), representing the boundary of the safe set in the $p_{x}-p_{y}$ plane with $v=0$ and $\Psi=-\pi/2$ (corresponding to the speed and heading of the rightmost initial state), evidences this explanation by showing that this initial state is outside the safe set.”
@@
In comparison, the 3DSF learned by RL refines the reference policy for all the initial states. The trajectory starting from the rightmost initial state in Fig. trajectory(b) demonstrates the superiority of the proposed 3DSF learned by RL. In particular, the safe APF policy makes the trajectory fail into a local optimum inside the right obstacle, while the unsafe APF policy with the 3DSF circumvents the local optimum and steers the system to the target.
@@
In Figs. trajectory(c-d), we show the trajectories of the vehicle controlled by the DDPG policies. It is found that although for some initial states safe DDPG successfully plans a safe trajectory, some trajectories fail to converge to the target set. This illustrates that it is hard to balance the task performance and constraint satisfaction by naively adding penalties to the stage cost during training. In contrast, the safety filter significantly enhances the safety performance of the unsafe DDPG policy while ensuring that the policy reaches the target.
@@
(a) Dashed blue curve: unsafe APF; Solid blue curve: unsafe APF with 3DSF learned from the expert controller; Red curve: safe APF.
@@
(b) Dashed blue curve: unsafe APF; Solid blue curve: unsafe APF with 3DSF learned by RL; Red curve: safe APF.
@@
(c) Dashed blue curve: unsafe DDPG; Solid blue curve: unsafe DDPG with 3DSF learned from the expert controller; Red curve: safe DDPG.
@@
(d) Dashed blue curve: unsafe DDPG; Solid blue curve: unsafe DDPG with 3DSF learned by RL; Red curve: safe DDPG.
@@
Figure 4: Comparison of closed-loop vehicle trajectories under different controllers and different safety filters. Black dots represent initial states. Green curve: the boundary of the safe set learned from the expert controller.
We have also implemented an MPC controller with a prediction horizon of 20 and a stage cost identical to that used in the DDPG controller. All state and input constraints are enforced as hard constraints, resulting in a constrained nonlinear optimization problem that is solved using active set. The MPC problem is warm-started using the shifted optimal solution from the previous time step. Fig. 5 illustrates the performance of the MPC controller, where the initial states are chosen to match those used for the other controllers in the simulation. As shown in the figure, all trajectories remain strictly within the feasible region. However, the trajectory starting from $p_{x}=2,p_{y}=0.5$ does not reach the target set. This is likely due to the non-convex nature of the MPC optimization problem, which may cause the solver to converge to a globally suboptimal control sequence.
Figure 5: Closed-loop trajectories of MPC.
@@ 8.4 Statistical evaluation and comparison with indirect data-driven safety filters
(c) Trajectories of the vehicle controlled by MPC.
Figure 6: Comparison of closed-loop vehicle trajectories under the 3DSF, the indirect safety filter, and MPC.
Finally, we statistically compare the proposed 3DSF against an indirect data-driven safety filter and MPC. In this subsection, the SACBF is learned by the RL approach because it has been illustrated in the previous subsection that the SACBF learned by the RL approach results in a larger safe set. To design the indirect data-driven safety filter, we intend to learn a standard CBF based on an approximate model. We first use a neural network with [128 128 128] “elu” hidden layers to identify the kinematic model. After 30 epochs of training, we get an approximate model with a sufficiently small ($2.38\cdot 10^{-4}$) mean square error. Then, following the approaches of [9, 17, 24], the approximate model is used as a prediction model in the reachability problem [^4] The reachability problem (4.1) is defined over an infinite horizon. To make it solvable for each sampled state, we truncate the horizon at 20. While a longer horizon provides a better approximation of the maximal safe set, a horizon of 20 balances between sample complexity and safety performance. (4.1). We solve (4.1) with the initial state $x_{0}$ being each state sample obtained in Section VIII.B. Then, we get the target value for the neural CBF, which is parameterized by a neural network with [128 128 32] “tansig” hidden layers.
Table 1: Comparison of safety rate, total cost, and average online CPU time for different control policies and safety filters.
@@
The results in Table 1 provide a detailed comparison of the safety rate, total cost, and average CPU time for those trajectories starting from the sampled initial states. Each controller in Table 1 is evaluated under three different settings: (1) No SF (no safety filter): The controller is executed directly without any safety mechanism; (2) Indirect SF: The controller is safeguarded by the indirect safety filter; (3) 3DSF: The controller is filtered by the proposed 3DSF. For each combination, we evaluate (i): Safety Rate: The percentage of simulated trajectories that satisfy all constraints; (ii) Total Cost: The accumulated performance cost ($p^{2}_{x}+p^{2}_{y}$ over 1000 steps), which represents the task performance; (iii) CPU Time: The average online computation time required for computing a control input. The key findings are summarized below:
@@
  • Although MPC enforces strict constraint satisfaction, its task performance is significantly worse than that of the other controllers (only 78.72% of trajectories eventually reach the target). This is probably due to the high non-convexity of the MPC optimization problem, which can lead to highly sub-optimal control inputs. In addition, the average CPU time of MPC is substantially higher than that of all other controllers.
@@ 9 Conclusions and future work
We have proposed an optimization-based control framework that incorporates an SACBF constraint to ensure safety for learning-based control methods. The key advantages of the framework are its compatibility with most RL and SL approaches for synthesizing safe controllers and its online computational efficiency. In particular, we demonstrate that any value-based RL algorithm or SL method can be used to construct both the objective function and the SACBF constraint in the optimization-based controller. Moreover, the method is online computationally efficient compared with other optimization-based control techniques that require long prediction horizons, such as MPC. Moreover, we have analyzed the theoretical properties regarding the robustness of SACBF and translated the results to error-to-state safety (ESSf) of the proposed control framework w.r.t. learning errors.
The limitations of our method include its sample complexity, the possibility of constraint violations during learning, and challenges in safety guarantees for the RL approach. Future work will focus on: (i) reducing sample complexity by leveraging model symmetries or reduced-order representations, (ii) considering continuous-time systems, and (iii) developing new formulations of value-function-type SACBFs that guarantee both safety and the convergence of RL algorithms.